OpenStudy (anonymous):

A regular deck of cards has 52 cards. There are four suits of 13 cards each. The suits are called clubs (black), diamonds (red), hearts (red), and spades (black). Two cards are drawn at random (without replacement) from a regular deck of cards. What is the probability that the first card is red and the second is red?

OpenStudy (anonymous):

(1/2)*(25/52)

OpenStudy (anonymous):

not 1/2 * 1/2 because when you choose the second card without replacing the first, there is one less red card to choose from.

OpenStudy (zarkon):

cjn566 is correct

OpenStudy (anonymous):

so 25/104?

OpenStudy (zarkon):

oops no he is not

OpenStudy (anonymous):

(1/2)*25/51

OpenStudy (zarkon):

\[\frac{26}{52}\cdot\frac{25}{51}\]

OpenStudy (zarkon):

didn't notice he had a 52 on the bottom

OpenStudy (anonymous):

the probablity is 25/102

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

thank you all!!!!! your really helping me understand this!

OpenStudy (anonymous):

25/102